# Randomly Sampled Language Reasoning Problems Reveal Limits of LLMs

### v2

@@ Abstract
Can LLMs pick up language structure from examples? Evidence in prior work seems to indicate yes, as pretrained models repeatedly demonstrate the ability to adapt to new language structures and vocabularies. However, this line of research typically considers languages that are present within common pretraining datasets, or otherwise share notable similarities with these seen languages. In contrast, in this work we attempt to measure models’ language understanding capacity while circumventing the risk of dataset recall. We parameterize large families of language tasks recognized by deterministic finite automata (DFAs), and can thus sample novel language reasoning problems to fairly evaulate LLMs regardless of training data. We find that, even in the strikingly simple setting of 3-state DFAs, LLMs underperform unparameterized $n$-Gram models on both language recognition and synthesis tasks. These results suggest that LLMs struggle to match the ability of basic language models in recognizing and reasoning over languages that are sufficiently distinct from the ones they see at training time, underscoring the distinction between learning individual languages and possessing a general theory of language.
@@ 1 Introduction
To circumvent this problem, we propose the following general approach: first define a large, exhaustive, and parsimoniously-defined space of languages that represents all languages of a certain difficulty level. Then, sample random languages from this space. By sampling randomly from this space, we can guarantee no bias towards canonical languages that might share structure with ones in the training dataset. In this work, we use languages recognized by 3-state DFAs as these are the lowest nontrivial difficulty level, but this technique can be generalized to produce benchmarks of any difficulty level.
Experimental results using this approach suggest that contemporary LLMs possess less sophisticated language pattern recognition abilities than expected; underperforming basic, parameter-free n-gram language models on even the simplest languages. These results, combined with LLMs’ impressive results on a variety of specific tasks, suggest that LLMs function as ensemble models over tasks they have seen in their dataset, but do not possess the ability to generalize to entirely novel language reasoning tasks.
@@ 2.1 Reasoning with LLMs
Reasoning is one of many “emergent abilities” (Wei et al., 2022a) possibly possessed by LLMs (Huang & Chang, 2022), although the nonlinear dependence of such emergent abilities on model size is disputed (Schaeffer et al., 2024). The chain-of-thought prompting technique (Wei et al., 2022b) has inspired a number of approaches to encourage latent reasoning ability of models (Yao et al., 2023; Besta et al., 2024; Kojima et al., 2022), including neuro-symbolic methods (Hua & Zhang, 2022; Weir et al., 2023, 2024). Building on this, other work considers how to optimize exemplars used for in-context learning (Dong et al., 2022) and chain-of-thought prompting, known as “rationale refinement” (Liu et al., 2021; Fu et al., 2022). Problem-decomposition is also shown to be effective (Zhou et al., 2022; Khot et al., 2022).
@@ 2.2 LLM reasoning evaluation
LLM reasoning abilities are often tested on natural language benchmarks and commonly seen problems like arithmetic (Cobbe et al., 2021; Amini et al., 2019; Hendrycks et al., 2021), commonsense reasoning (Bhargava & Ng, 2022), and other, sometimes generative, tasks (Lake & Baroni, 2018; Pasupat & Liang, 2015; Lin et al., 2019) and task collections (Srivastava et al., 2022). LLMs have been shown to lack sufficient reasoning capability across a range of tasks including multi-step planning and complex inference (Valmeekam et al., 2022). Fan et al. (2023) introduce an LLM reasoning benchmark on algorithmic problems through NP-hard complexity, and Hazra et al. (2024) show that LLMs struggle to complete simple 3SAT problems. Patel et al. (2021) demonstrate that much of LLM mathematical reasoning can be explained by shallow heuristics, and Razeghi et al. (2022) similarly find that term frequency in training data impacts models’ in-context learning ability. In comparison to these, we attempt to explore the distinction described by Patel et al. (2021), but push both language simplicity and language unfamiliarity to their limits, by exploring simple languages recognized by randomly sampled DFAs. This enables us to evaluate the ability of LLMs to reason about language.
@@ 3.1 DFAs and Regular Languages
The original Chomsky Hierarchy (Chomsky, 1959) separates language into four types (Figure 2). We focus on the task of understanding Type 3 languages, the simplest form of language in the hierarchy, that are recognized by a Deterministic Finite Automaton (DFA). Examples of languages recognized by DFAs include simple examples like binary strings with an even number of ones, and even such examples as numbers in base 10 divisible by 7. Finite automaton languages are equivalent to regular languages, which are recognized by regular expressions.
One simple metric of the difficulty of a regular language is the number of states in the corresponding DFA, which represents the amount of memory the automaton has at any point while processing a given sequence. [^1] There are other metrics of difficulty, but we choose number of states as it is highly parsimonious. DFAs with 2 states have the property that their set of states is isomorphic to the output set $\{0,1\}$, and, therefore, do not have any hidden state. We thus explore 3-state DFAs, as this is the simplest nontrivial case.
@@ 3.2 Sequence Completion Task
We also pose a sequence completion task, in which models must complete a sequence in a given DFA’s language. We study this task primarily because, in practice, most language data that models encounter will be in roughly this format, with several example sequences in a given language followed by a distinct prefix that needs to be completed via next token prediction.
To generate test cases for this task, we (1) sample a DFA, (2) sample 30 example sequences of length 10 that this DFA accepts, and (3) sample a distinct prefix of length 5 that is not a prefix of any of our 30 example sequences, with the property that there exists some length-$\leq 5$ completion of this prefix that the DFA would accept. The task is to find a completion [^2] Not necessarily the same completion found in sampling. of this prefix of between 1 and 5 characters such that the DFA accepts the full sequence. For details on sampling, see Appendix A.2.
We evaluate models by (1) sampling a DFA, (2) sampling 30 problem instances at random, and then (3) computing a binary prediction score (whether or not the predicted completion creates a valid string in the language) for each instance separately. We then average this metric over several sampled DFAs to produce a single accuracy score.
@@ 3.3 Transducer Task
While the sequence completion task is the natural one that comes to mind as a basic language task, it has a difficulty-gap problem. Specifically, the issue is that many DFAs, including the one shown in Figure 1, recognize languages that are particularly difficult to identify based on a set of examples, unless you build some kind of world model. Other DFAs end up being trivial to generate a completion for by analyzing common suffixes.333 The difficulty gap exists because a set of recognized sequences of length 10 gives no direct insight into intermediate states between the first and tenth token. As such, to be able to utilize this information for languages like the one in Figure 1 where there are no “resets” (sequences of symbols that necessarily lead to a particular state), a model must be capable of hollistically evaluating the entire sequence, probably requiring a world model. Many other DFAs contain these resets, but do so in such a way that makes it possible to e.g., recognize that all sequences in this language end in a, making the problem trivial. To provide a more direct evaluation of non-world-modeling-based pattern recognition, we explore the Transducer task.
@@ 3.4 Baselines
l baseline models with varying degrees of sophistication.
@@
  • Brute-ForceS: take all possible DFAs with 3 states and 3 symbols. Filter for ones that accept all the example sequences. Then try all remaining DFAs on all $3^{5}$ possible 5-length completions and return the completion that the maximal number of DFAs accept.
@@
  • NullT baseline: for a given DFA, whichever of the following strategies produces a better result: always predict 0 or always predict 1.
@@ 4 Experiments
Finally, we evaluated the proprietary models GPT-3.5-turbo-instruct, GPT-3.5 Chat (turbo-0125) (OpenAI, 2024a), GPT-4o-mini (2024-07-18), GPT 4o (2024-05-13) (OpenAI, 2024b), and Claude 3.5 Sonnet (Anthropic, 2024).
For each open source model, we used a local VLLM (Kwon et al., 2023) server for evaluation and always evaluated on 1000 distinct DFAs. For GPT-4o and Claude, we evaluated on 30 DFAs due to computation costs. For all other proprietary models, we evaluated on 100 DFAs. All results are reported with 95% bootstrap confidence intervals.
@@
  • Basic provides no context, presenting the problem as a generic next-token prediction or sequence generation task, where output is provided immediately following the input, with no space to think.
  • More-Expl provides some general information about the grammar that the input strings are generated from. This remains a next token prediction/sequence generation task.
  • COT provides general information about the grammar, and additionally invokes chain-of-thought reasoning to help the model reason over the task. Here, the model is given space to reason before providing a tagged answer.
@@
We produce versions of each of these prompts for each task, denoting these with a subscript S for sequence completion prompts and T for transducer prompts. Full listings of these prompts can be found in Appendix D, with Table 4 containing a summary of each prompt. While no finite set of prompts will be fully sufficient to capture all possible model behavior, we believe our set of prompts captures common prompting strategies.
@@ 5 Results
Main results for all tasks are presented in Table 1. For all LLMs, we ignore non-answers, i.e., if for a given DFA a model gets 25 correct answers, 1 incorrect answer, and responds with an unparseable result on 4, this counts as a 25/26, not a 25/29.
@@ 5.1 Sequence Completion
As seen in Table 1, this task is nearly always fully determined, that is, it can be solved with $\sim$100accuracy in theory, as demonstrated by BruteForceS results. Of course, BruteForceS is extremely computationally expensive, and, as such, we primarily focus on the $n$-GramS heuristics as our baselines. Still, we find, similarly to the transducer task, that $n$-GramS heuristics tend to outperform LLMs.
@@
One potential problem with using this task for cross-model comparisons is the relevance of tokenization. Unfortunately, we found that forcing uniform tokenization by using commas in the prompt uniformly reduced accuracy, see Appendix B for details.
@@
Table 2: Results for chat models (where we investigated multiple prompts). For all other models, we only used the Basic prompt. We find that non-COT prompts consistently work better for the Transducer task, with more mixed results on sequence completion.
@@ 5.2 Transducer
Unlike sequence completion, this task is generally not fully determined, with the $\textsc{BruteForce}_{T}$ model only achieving 96.4% accuracy. However, relative results should still be valid as the impossible problems are present with equal probability for all models.
We find that in general all LLMs underperform a 4-GramT model, demonstrating that they are unable to adequately solve this task. The relative performance of the models also does not generally correspond to their overall scale, with open source LLama-3 and Mistral Nemo 8B parameter models outperforming Claude and GPT-4o. Even within a model class we find no clear pattern: GPT-4o is outperformed by GPT 3.5, Llama 3-70B has similar performance to Llama 3-8B, and the Mistral Nemo 12B models perform similarly to Nemo Minitron 8B. Coding models also demonstrate no particular advantage on this task.
The generally lower performance of chat-oriented models suggests this task is better suited to non-chat models. To investigate that this is not specific to the Basic prompt, we investigate several chat models on the other prompts. As seen in Table 2, our chain-of-though and word problem prompts, which attempt to leverage the full reasoning capabilities of chat models, also fare poorly, performing similarly or worse to the Basic prompt on the Transducer task.
@@ 5.3 Case Study: Sum Modulo 3 DFA
$>3$-Gram 4 5
Sum Modulo 3 0 0
Figure 3: Results on Sum Modulo 3 DFA. (a) Venn diagram results for CR (claude-3.5 using Red-GreenT), QB (qwen-2.5-coder-instruct-7B using BasicT), and 6-GramT. Percentages are accuracy numbers, and venn diagram is error counts, out of 1000. (b) Results of qualitative analysis. 30 CR reasoning traces were sampled in each condition, with Incorrect traces guaranteed to correspond to problems 6-GramT got correct.
We focus on the sum modulo 3 DFA in this section. We depict this DFA in Figure 1. This DFA can be interpreted as an arithmetic check, where a represents 0, b represents 1, and c represents 2, and the DFA accepts strings whose sum is equal to 0 modulo 3. We focus on the Transducer setting as sum modulo 3 is nearly impossible (for LLMs and $n$-Gram models alike) in the sequence completion setting due to the issues raised in Section 3.3. For this case study, we focus on the model/prompt combinations QB and CR, defined as
  • QB: qwen-2.5-coder-instruct-7B with BasicT. Selected as it is the best performing combination overall.
  • CR: claude-3.5 with Red-GreenT. Selected as it is the best performing combination that provides an explanation (needed later for our qualitative analysis)
@@
Despite transformers’ theoretically high computational capacity, without the ability to pattern match to existing problems, Claude uses an unsophisticated and ineffectual approach.
@@ 6 Conclusion
These results expose a particular weakness in LLMs, as many potential foundation model applications involve tasks that are not expressed in familiar human languages or pre-existing programming languages. More specifically, in tasks where there is a need to produce an output in a precise, atypical, format, we should be skeptical of the ability of LLMs to in-context-learn this format. For these tasks, it may be prudent to seek a new approach.
Acknowledgments
We thank Marc Marone and Theo Olausson for their advice on experiment design.

### v3

@@ Abstract
Can LLMs pick up language structure from examples? Evidence in prior work seems to indicate yes, as pretrained models repeatedly demonstrate the ability to adapt to new language structures and vocabularies. However, this line of research typically considers languages that are present within common pretraining datasets, or otherwise share notable similarities with seen languages. In contrast, in this work we attempt to measure models’ language understanding capacity while circumventing the risk of dataset recall. We parameterize large families of language tasks recognized by deterministic finite automata (DFAs), and can thus sample novel language reasoning problems to fairly evaluate LLMs regardless of training data. We find that, even in the strikingly simple setting of 3-state DFAs, LLMs underperform unparameterized $n$-Gram models on both language recognition and synthesis tasks. These results suggest that LLMs struggle to match the ability of basic language models in recognizing and reasoning over languages that are sufficiently distinct from the ones seen at training time, underscoring the distinction between learning individual languages and possessing a general theory of language.
@@ 1 Introduction
To circumvent this problem, we propose the following general approach: first define a large, exhaustive, and parsimoniously-defined space of languages that represents all languages of a certain difficulty level. Then, sample random languages from this space. By sampling randomly, we can guarantee no bias towards canonical languages that might share structure with ones in the training dataset. In this work, we use languages recognized by 3-state DFAs as these are the lowest nontrivial difficulty level, but this technique can be generalized to produce benchmarks of any difficulty level.
Experimental results using this approach suggest that contemporary LLMs possess less sophisticated language pattern recognition abilities than expected; underperforming basic, parameter-free n-gram language models on even the simplest languages. These results, combined with LLMs’ impressive results on a variety of specific tasks, suggest that LLMs function as ensemble models over language tasks they have seen in their dataset, but do not possess the ability to generalize to entirely novel language reasoning tasks.
@@ 2.1 Reasoning with LLMs
Reasoning is one of many “emergent abilities” (Wei et al., 2022a) possibly possessed by LLMs (Huang & Chang, 2022), although the nonlinear dependence of such emergent abilities on model size is disputed (Schaeffer et al., 2024). The chain-of-thought prompting technique (Wei et al., 2022b) has inspired a number of approaches to encourage the latent reasoning ability of models (Yao et al., 2023; Besta et al., 2024; Kojima et al., 2022), including neuro-symbolic methods (Hua & Zhang, 2022; Weir et al., 2023, 2024). Building on this, other work considers how to optimize exemplars used for in-context learning (Dong et al., 2022) and chain-of-thought prompting, known as “rationale refinement” (Liu et al., 2021; Fu et al., 2022). Problem-decomposition is also shown to be effective (Zhou et al., 2022; Khot et al., 2022).
@@ 2.2 LLM reasoning evaluation
LLM reasoning abilities are often tested on natural language benchmarks and commonly seen problems like arithmetic (Cobbe et al., 2021; Amini et al., 2019; Hendrycks et al., 2021), commonsense reasoning (Bhargava & Ng, 2022), and other, sometimes generative, tasks (Lake & Baroni, 2018; Pasupat & Liang, 2015; Lin et al., 2019) and task collections (Srivastava et al., 2022). LLMs have been shown to lack sufficient reasoning capability across a range of tasks including multi-step planning and complex inference (Valmeekam et al., 2022). Fan et al. (2023) introduce an LLM reasoning benchmark on algorithmic problems through NP-hard complexity, and Hazra et al. (2024) show that LLMs struggle to complete simple 3SAT problems. Patel et al. (2021) demonstrate that much of LLM mathematical reasoning can be explained by shallow heuristics, and Razeghi et al. (2022) similarly find that term frequency in training data impacts models’ in-context learning ability. In comparison to these, we explore the distinction described by Patel et al. (2021), but push both language simplicity and language unfamiliarity to their limits, by exploring simple languages recognized by randomly sampled DFAs. This enables us to evaluate the ability of LLMs to reason about language.
@@ 3.1 DFAs and Regular Languages
The original Chomsky Hierarchy (Chomsky, 1959) separates language into four types (Figure 2). We focus on the task of understanding Type 3 languages, the simplest form of language in the hierarchy, that are recognized by a Deterministic Finite Automaton (DFA). Examples of languages recognized by DFAs include simple ones like binary strings with an even number of ones, and even such examples as numbers in base 10 divisible by 7. Finite automaton languages are equivalent to regular languages, which are recognized by regular expressions.
One simple metric of the difficulty of a regular language is the number of states in the corresponding DFA, which represents the amount of memory the automaton has at any point while processing a given sequence. [^1] There are other metrics of difficulty, but we choose number of states as it is highly parsimonious. DFAs with 2 states have the property that their set of states is no larger than the output set $\{0,1\}$, and, therefore, do not have any hidden state. We thus explore 3-state DFAs, as this is the simplest nontrivial case.
@@ 3.2 Sequence Completion Task
We first pose a sequence completion task, in which models must complete a sequence in a given DFA’s language. We study this task primarily because, in practice, most language data that models encounter will be in roughly this format, with several example sequences in a given language followed by a distinct prefix that needs to be completed via next token prediction.
To generate test cases for this task given a DFA, we first (1) sample 30 example sequences of length 10 that this DFA accepts, and then (2) sample a distinct prefix of length 5 that is not a prefix of any of our 30 example sequences, with the property that there exists some length-$\leq 5$ completion of this prefix that the DFA would accept. The task is to find a completion (not necessarily the same completion found in sampling) of this prefix of between 1 and 5 characters such that the DFA accepts the full sequence. For details on sampling, see Appendix A.2.
We evaluate models by (1) sampling a DFA, (2) sampling 30 problem instances at random (each of which contains 30 example sequences and a distinct prefix), and then (3) computing a binary prediction score (whether or not the predicted completion creates a valid string in the language) for each instance separately. We then average this metric over several sampled DFAs to produce our accuracy score.
@@ 3.3 Transducer Task
While the sequence completion task is the natural one that comes to mind as a basic language task, it has a difficulty-gap problem. Specifically, the issue is that many DFAs, including the one shown in Figure 1, recognize languages that are particularly difficult to identify based on a set of examples, unless you build some kind of world model. Other DFAs end up being trivial to generate a completion for by analyzing common suffixes.222 The difficulty gap exists because a set of recognized sequences of length 10 gives no direct insight into intermediate states between the first and tenth token. As such, to be able to utilize this information for languages like the one in Figure 1 where there are no “resets” (sequences of symbols that necessarily lead to a particular state), a model must be capable of hollistically evaluating the entire sequence, probably requiring a world model. Many other DFAs contain these resets, but do so in such a way that makes it possible to e.g., recognize that all sequences that end in a are in the language, making the problem trivial. To provide a more direct evaluation of non-world-modeling-based pattern recognition, we explore the Transducer task.
@@ 3.4 Baselines
l To contextualize LLM accuracies, waseline models with varying degrees of sophistication.
@@
  • Brute-ForceS: take all possible DFAs with 3 states and 3 symbols. Filter for ones that accept all the example sequences. Then try all remaining DFAs on all $3^{5}$ possible 5-length completions and return the completion that the maximal number of DFAs accept, breaking ties arbitrarily.
@@
  • NullT baseline: for a given DFA, whichever of the following strategies produces a higher accuracy: always predict 0 or always predict 1.
@@ 4 Experiments
Finally, we evaluated the proprietary models GPT-3.5-turbo-instruct, GPT-3.5 Chat (turbo-0125) (OpenAI, 2024a), GPT-4o-mini (2024-07-18), GPT 4o (2024-05-13) (OpenAI, 2024b), o1-preview (2024-09-12) (OpenAI et al., 2024), and Claude 3.5 Sonnet (Anthropic, 2024).
For each open source model, we used a local VLLM (Kwon et al., 2023) server for evaluation and always evaluated on 1000 distinct DFAs. For GPT-4o and Claude, we evaluated on 30 DFAs due to computation costs. For o1-preview we evaluated on only 10 DFAs, and only on the Transducer task (which we felt was a better fit for a reasoning model). For gpt-3.5 and gpt-4o-mini, we evaluated on 100 DFAs. All models were evaluated with temperature 0, except o1-preview333o1-preview does not allow setting a non-default temperature..
@@
  • Basic provides no context, presenting the problem as a generic sequence generation or next-token prediction task, where output is provided immediately following the input, with no space to think.
  • More-Expl explains that the strings are generated from a DFA with 3 states, but is otherwise identical to Basic. This remains a sequence generation/next token prediction task.
  • COT provides the same information as More-Expl and additionally invokes chain-of-thought reasoning to help the model reason over the task. Here, the model is given space to reason before providing a tagged answer.
@@
We produce versions of each of these prompts for each task, denoting these with a subscript S for sequence completion prompts and T for transducer prompts. Full listings of these prompts can be found in Appendix E, with Table 5 containing a summary of each prompt. While no finite set of prompts will be fully sufficient to capture all possible model behavior, we believe our set of prompts captures common prompting strategies.
@@ 5 Results
Main results for all tasks are presented in Table 1. For all LLMs, we ignore non-answers, i.e., if for a given DFA a model gets 25 correct answers, 1 incorrect answer, and responds with an unparseable result on 4, this counts as a 25/26, not a 25/29. We then report the mean across DFAs and 95% bootstrap confidence intervals.
@@ 5.1 Sequence Completion
As seen in Table 1, this task is nearly always fully determined, that is, it can be solved with $\sim$100accuracy in theory, as demonstrated by BruteForceS results. Of course, BruteForceS is extremely computationally expensive, and, as such, we primarily focus on the $n$-GramS heuristics as our baselines. Still, we find that $n$-GramS heuristics tend to outperform LLMs.
@@
One potential problem with using this task for cross-model comparisons is the relevance of tokenization. Unfortunately, we found that forcing uniform tokenization by using commas in the prompt uniformly reduced accuracy, see Appendix D for details.
@@
Table 2: Results for models where we investigated multiple prompts (we only used Basic on other models). We bold the best prompt for each model. Non-COT prompts consistently work better for the Transducer task, with more mixed results on sequence completion.
@@ 5.2 Transducer
Unlike sequence completion, this task is generally not fully determined, with the $\textsc{BruteForce}_{T}$ model only achieving 96.4% accuracy. However, relative results should still be valid as the impossible problem instances are present with equal probability for all models.
We find that in general all LLMs underperform a 4-GramT model, demonstrating that they are unable to adequately solve this task. The relative performance of the models also does not correspond to their overall scale, with open source LLama-3 and Mistral Nemo 8B parameter models outperforming Claude and GPT-4o. Even within a model class we find no clear pattern: GPT-4o and o1-preview [^4] The particularly poor performance of o1-preview may be due to the model not supporting temperature 0. In other experiments, we found that GPT models with temperature 1 tended to perform poorly. See Appendix B for more details. are outperformed by GPT 3.5, Llama 3-70B has similar performance to Llama 3-8B, and the Mistral Nemo 12B models perform similarly to Nemo Minitron 8B. Coding models also demonstrate no advantage on this task.
The generally lower performance of chat-oriented models suggests this task is better suited to non-chat models. To investigate that this is not specific to the Basic prompt, we investigate other prompts for chat models. As seen in Table 2, our chain-of-though and word problem prompts, which attempt to leverage the full reasoning capabilities of chat models, also fare poorly, performing similarly or worse to the Basic prompt on the Transducer task in all cases.
@@
5.3 Comparison of Benchmarks
Figure 3: Transducer and sequence completion results plotted against each other. Points represent the mean over several DFAs and intervals represent 95% confidence intervals. Points are colored by model type, with the best and worst model by each metric in each category labeled, as well as all baseline and proprietary models.
Figure 3 displays the relationship between model performance on the Sequence Completion and Transducer benchmarks. While at a high level, there is a positive correlation between the two, there are a few notable differences. For one, the Code models perform notably better than other open source models on Sequence Completion, but not on Transducer. Additionally, on Transducer, a ceiling on performance is observed, where LLMs cluster together between 3-GramT and 4-GramT performance; this clustering does not appear on the Sequence Completion benchmark.
@@ 5.4 Case Study: Sum Modulo 3 DFA
2 red rooms 2 3
Figure 4: Results on Sum Modulo 3 DFA. (a) Venn diagram of errors (out of 1000). Labeled percentages are accuracies. (b) Results of qualitative analysis, out of 30 in both cases.
We investigate the transducer task on the DFA depicted in Figure 1. This DFA can be interpreted as an arithmetic check, where a represents 0, b represents 1, and c represents 2, and the DFA accepts strings whose sum is equal to 0 modulo 3. For this case study, we focus on the model/prompt combinations MB and CR, defined as
  • MB: mistral-nemo-minitron-8B/BasicT. Selected as it is the best performing combination overall.
  • CR: claude-3.5/Red-GreenT. Selected as it is the best performing combination that provides an explanation (needed later for our qualitative analysis)
@@
Despite transformers’ high computational capacity, without the ability to pattern match to existing problems, Claude uses an unsophisticated and ineffectual approach.
@@ 6 Conclusion
Many potential foundation model applications involve tasks that are not expressed in familiar human languages or pre-existing programming languages. More specifically, in tasks where there is a need to produce an output in a precise, atypical, format, we should be skeptical of the ability of LLMs to in-context-learn this format. For these tasks, it may be prudent to seek a new approach.
Impact Statement
Aside from the social consequences of this work as related to advancing the field of Machine Learning in general, this work has the goal of advancing the field of benchmarks in Machine Learning. While we view this as a positive objective, as it ensures that models are being evaluated fairly, it might have negative consequences insofar as benchmarking techniques might be best left unpublished to prevent deliberate or unintentional overfitting.